# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v3

@@ Abstract
Electronic Health Record (EHR) data, while rich in information, often suffers from sparsity, posing significant challenges in predictive modeling. Traditional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies in models. Addressing this, we introduce PRISM, a novel approach that indirectly imputes data through prototype representations of similar patients, thus ensuring denser and more accurate embeddings. PRISM innovates further with a feature confidence learner module, which evaluates the reliability of each feature in light of missing data. Additionally, it incorporates a novel patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III and MIMIC-IV datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
footnotetext: $\dagger$ Corresponding author.
@@ 1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have garnered significant attention in leveraging EHR data to inform clinical decisions. Such applications range from predicting the survival risk of patients Ma et al. (2020b); Zhang et al. (2021) to forecasting early mortality outcomes Yan et al. (2020); Ma et al. (2022); Gao et al. (2023); Zhu et al. (2023a).
@@
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption Chen and Zhang (2020), expensive examinations Ford and Ford (2000), and safety considerations Zhou et al. (2020) result in missing observations; for instance, not all indicators are captured during every patient visit Che et al. (2018). Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy Zhang et al. (2022). Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE Van Buuren and Groothuis-Oudshoorn (2011), non-negative matrix factorization (NMF) Wang and Zhang (2012), and compressed sensing Lopes (2013). However, they fail to capture the time-series characteristics of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance Ma et al. (2020a, 2023) exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies Liu et al. (2023). Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases Suo et al. (2018). Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data Zhang et al. (2021, 2022). Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity Lee et al. (2015), handcrafted metrics Huang et al. (2019), or learning-based methods Suo et al. (2018), interpret these values identically. This results in a potentially misleading perception of similarity.
@@
In healthcare, the absence of data can severely compromise confidence in a prognosis. Addressing and understanding the challenges of these missing features is of utmost importance. In light of this, our work, PRISM seeks to bridge this gap. Our primary contributions are as follows:
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Such feature-missing awareness through the whole pipeline allows for a more refined patients representation.
  • Experimentally, comprehensive experiments on two real-world datasets, MIMIC-III and MIMIC-IV, focusing on in-hospital mortality and 30-day readmission prediction tasks, reveal that PRISM significantly improves the quality of patient representations against EHR data sparsity. It outperforms the best-performing baseline model with relative improvements of 6.40% and 2.78% in mortality and 1.38% and 1.63% in readmission on AUPRC for MIMIC-III and MIMIC-IV, respectively. Further ablation studies and detailed experimental analysis underline PRISM’s effectiveness, robustness, adaptability, and efficiency.
@@ 2 Related Work
In the realm of EHR data analysis, irregular sampling often leads to significant data sparsity, presenting substantial challenges in modeling. Previous methods on sparse EHR data predominantly fall under two categories: indirect and direct imputation.
@@
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE Van Buuren and Groothuis-Oudshoorn (2011), non-negative matrix factorization (NMF) Wang and Zhang (2012), compressed sensing Lopes (2013), or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D Che et al. (2018) adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare Ma et al. (2020b) and AICare Ma et al. (2023) first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
@@
An alternative approach is evident in GRASP Zhang et al. (2021) and M3Care Zhang et al. (2022), which emphasizes the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis Zhang et al. (2021). However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  Lee et al. (2015) and  Huang et al. (2019), have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  Suo et al. (2018), which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
@@ 3 Problem Formulation
EHR datasets consist of a sequence of dynamic and static information for each patient. Assuming that there are $F$ features in total, $D$ dynamic features (e.g., lab tests and vital signs) and $S$ static features (e.g., sex and age), where $F=D+S$, at every clinical visit $t$. The features recorded at visit $t$ can be denoted as $\bm{x}_{t}\in\mathbb{R}^{F},t=1,2,\cdots,T$, with total $T$ visits. The dynamic feature information can be formulated as a 2-dimensional matrix $\bm{d}\in\mathbb{R}^{T\times D}$, along with static information denoted as 1-dimensional matrix $\bm{s}\in\mathbb{R}^{S}$. In addition, to differentiate between categorical and numerical variables within dynamic features, we employ one-hot encoding for categorical variables. Due to the inherent sparsity of EHR data, we incorporate feature missingness as part of our inputs. At a global view, we define the missing representation, denoted as $\rho_{i}$, to be the presence rate of the $i$-th feature within the entire dataset. From a local view, the missing representation, $\tau_{i,t}$, signifies the time interval since the last recorded visit that contains the $i$-th feature up to the $t$-th visit.
@@
Prediction objective is presented as a binary classification task. Given each patient’s EHR data $\bm{X}=[\bm{x_{1}},\bm{x_{2}},\cdots,\bm{x_{T}}]^{\top}\in\mathbb{R}^{T\times F}$ and feature missing status $\{\rho,\tau\}$ as input, where each $\bm{x_{t}}$ consists of dynamic features and static features representation, the model attempts to predict the specific clinical outcome, denoted as $y$. The objective is formulated as $\hat{y}=\text{Model}(\bm{X},\{\rho,\tau\})$. In the MIMIC dataset’s in-hospital mortality prediction task, the goal is to predict the discharge status (00 for alive, $1$ for deceased) based on the initial 48-hour window of an ICU stay. Similarly, the 30-day readmission task predicts if a patient will be readmitted in 30 days (00 for no readmission, $1$ for readmission).
@@ 4.1 Overview
Figure 1: Overall model architecture of our proposed method. “Rep.” means “Representation”.
@@
  • Feature-Isolated Embedding Module applies GRU and MLP backbone separately to dynamic features and static features. Each dynamic feature learns historical representations over multiple time steps. In order to align with the original attribute information of each feature, the features are learned in isolation from each other.
@@
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies GNN to learn prototype similar patients.
@@ 4.2 Feature-Isolated Embedding
In this module, static and dynamic features are learned individually, yielding feature representations of unified dimensions $f$.
Static Features Embedding Static features remain constant at each visit. Hence, we opt for a single-layer MLP to map each static feature into the feature dimensions $f$:
@@
Dynamic Features Embedding To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU:
@@ 4.3 Feature Confidence Learner
Existing models often adopt various ways to enhance patient representations in order to mitigate the noise introduced by processing sparse EHR data. However, these models often use imputed data and ignore the impact of feature missing status, thus reducing the credibility of the learned hidden representation. We design a measurement called “feature confidence”, which represents the reliability of the input feature values for each patient and each visit. In addition, we have incorporated this measure into the self-attention mechanism as a recalibration module to elevate low-confidence features’ attention.
@@ 4.4 Confidence-Aware Prototype Patient Learner
To comprehensively account for the implications of absent features in the evaluation and identification of analogous patients, we integrate the feature confidence status into our confidence-aware module for discovering cohorts of similar patients. Based on the graph convolutional network (GCN), we compute the similarity score by factoring in the acquired feature confidence pertaining to the individual patient. This approach facilitates the preferential selection of patients who are more similar to the subject patient.
@@
Prototype Patients Cohort Discovery To compute the enhanced representation information of similar patients, we design the prototype similar patient cohort discovery module. First, we utilize the K-Means clustering algorithm to cluster all $F$ features into $K$ groups. Then we select the center vectors of $K$ clusters as the initial $K$ prototypes $\ \bm{\mathcal{G}_{k}},\ k=1,2,\cdots,K$. With convolutions occurring within these graphs, the graph structure dynamically evolves across epochs.
Based on the aforementioned similarity measure, we can find the most similar prototype for each sample, making samples within the clusters more similar to each other during the learning process; at the same time, the prototypes also become more representative during the learning process.
After concatenating the hidden representation $\bm{z}$, computed in the previous step, with $\bm{\mathcal{G}}$, we feed this along with the similarity matrix $\bm{A}$ into a GCN. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. However, during training, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adjusted. Consequently, we use the respective $\bm{\mathcal{G}}$ for each patient for representation enhancement:

### v4

@@ Abstract
Electronic Health Record (EHR) data, while rich in information, often suffers from sparsity, posing significant challenges in predictive modeling. Traditional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies in models. Addressing this, we introduce PRISM, a framework that indirectly imputes data through prototype representations of similar patients, thus ensuring denser and more accurate embeddings. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature in light of missing data. Additionally, it incorporates a new patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
footnotetext: Corresponding author.
@@ 1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have garnered significant attention in leveraging EHR data to inform clinical decisions. Such applications range from predicting the survival risk of patients [1, 2] to forecasting early mortality outcomes [3, 4, 5, 6].
@@
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption [7], expensive examinations [8], and safety considerations [9] result in missing observations; for instance, not all indicators are captured during every patient visit [10]. Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy [11]. Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE [12], non-negative matrix factorization (NMF) [13], and compressed sensing [14]. However, they fail to capture the time-series characteristics of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance [15, 16] exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies [17]. Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases [18]. Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data [2, 11]. Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity [19], handcrafted metrics [20], or learning-based methods [18], interpret these values identically. This results in a potentially misleading perception of similarity.
@@
In healthcare, the absence of data can severely compromise confidence in a prognosis. Addressing and understanding the challenges of these missing features is of utmost importance. In light of this, PRISM seeks to bridge this gap. Our primary contributions are:
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Compared to existing SOTA baselines, i.e. GRASP [2] and M3Care [11], PRISM provides a refined feature calibration, and further missing-aware similarity measure helps to identify more related patient representation, with missing feature status elaborately taking into account.
  • Experimentally, comprehensive experiments on four real-world datasets, MIMIC-III, MIMIC-IV, Challenge-2012, and eICU, focusing on in-hospital mortality and 30-day readmission prediction tasks, reveal that PRISM significantly improves the quality of patient representations against EHR data sparsity. PRISM outperforms the best-performing baseline model with relative improvements of 6.40%, 2.78%, 1.51% and 11.01% in mortality on AUPRC for four datasets. In terms of readmission task, PRISM obtains relative improvements of 1.38% and 1.63% on AUPRC for MIMIC-III and MIMIC-IV, respectively. Further ablation studies and detailed experimental analysis underline PRISM’s effectiveness, robustness, adaptability, and efficiency.
@@ 2 Related Work
In the realm of EHR data analysis, irregular sampling often leads to significant data sparsity, presenting substantial challenges in modeling. Previous methods on sparse EHR data predominantly fall under two categories: direct imputation in the raw data space and indirect imputation within the feature representation space.
@@
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE [12], non-negative matrix factorization (NMF) [13], compressed sensing [14], or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D [10] adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare [1] and AICare [16] first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
@@
As demonstrated in GRASP [2] and M3Care [11], they emphasize the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis [2]. However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  [19] and  [20], have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  [18], which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
@@ 3 Problem Formulation
Table 1: Notations symbols and their descriptions
Notations Descriptions
$N$ Number of patient samples
$T$ Number of visits for a certain patient
$D$ Number of dynamic features
$S$ Number of static features
$F$ Number of features, $F=S+D$
$\bm{d_{it}}\in\mathbb{R}^{m}$ The $i$-th feature at the $t$-th visit, where $m$ is either the number of categories (for one-hot encoding) or 1 (for numerical lab tests)
$\bm{X}\in\mathbb{R}^{T\times F}$ Clinical visit matrix of a single patient, consisting of $T$ visits
$\bm{s},\bm{d}$ Static and dynamic feature vector of a patient
$\bm{y},\bm{\hat{y}}$ Ground truth labels and prediction results
$\bm{h_{i}}\in\mathbb{R}^{T\times f},\bm{h}$ Representation of $i$-th feature learned by GRU (for dynamic features) or MLP (for static features), stacked to form the representation matrix $\bm{h}$, $f$ is each feature’s embedding dimension
$\rho_{i}$ Feature presence rate of feature $i$ in training set
$\tau_{i,t}$ Time interval from the last recorded visit of $i$-th feature at $t$-th visit
$\bm{C}\in\mathbb{R}^{T\times F}$ Learned feature confidence matrix of a patient
$\bm{z_{i}}$ $\bm{z_{i}}$ is the learned representation of the $i$-th patient after the feature calibration layer
$\bm{\alpha},\bm{\alpha^{*}}$ Learned attention weights and calibrated attention weights after the feature calibration layer
$\phi(\cdot,\cdot)$ Patient similarity measure function
$\bm{A}=(a_{i,j})$ Adjacency matrix of patients, composed of similarity score between $i$-th and $j$-th patient
$\bm{\mathcal{G}_{k}}$ Learned prototype patient representation of the $k$-th group
$\bm{z^{*}_{i}}$ Learned representation of the $i$-th patient after representation fusion layer
$\bm{W}_{\square}$ Parameter matrices of linear layers. Footnote $\square$ denotes the name of the layer
$K$ Number of similar patient groups
@@
EHR datasets consist of a sequence of dynamic and static information for each patient. Assuming that there are $F$ features in total, $D$ dynamic features (e.g., lab tests and vital signs) and $S$ static features (e.g., sex and age), where $F=D+S$, at every clinical visit $t$. The features recorded at visit $t$ can be denoted as $\bm{x}_{t}\in\mathbb{R}^{F},t=1,2,\cdots,T$, with total $T$ visits. The dynamic feature information can be formulated as a 2-dimensional matrix $\bm{d}\in\mathbb{R}^{T\times D}$, along with static information denoted as 1-dimensional matrix $\bm{s}\in\mathbb{R}^{S}$. In addition, to differentiate between categorical and numerical variables within dynamic features, we employ one-hot encoding for categorical variables. Due to the inherent sparsity of EHR data, we incorporate feature missingness as inputs. At a global view, we define the missing representation, denoted as $\rho_{i}$, to be the presence rate of the $i$-th feature within the entire dataset. From a local view, the missing representation, $\tau_{i,t}$, signifies the time interval since the last recorded visit that contains the $i$-th feature up to the $t$-th visit.
@@
Prediction objective is presented as a binary classification task. Given each patient’s EHR data $\bm{X}=[\bm{x_{1}},\bm{x_{2}},\cdots,\bm{x_{T}}]^{\top}\in\mathbb{R}^{T\times F}$ and feature missing status $\{\rho,\tau\}$ as input, where each $\bm{x_{t}}$ consists of dynamic features and static features representation, the model attempts to predict the specific clinical outcome, denoted as $y$. The objective is formulated as $\hat{y}=\text{Model}(\bm{X},\{\rho,\tau\})$. For the in-hospital mortality prediction task, the goal is to predict the discharge status (00 for alive, $1$ for deceased) based on the initial 48-hour window of an ICU stay. Similarly, the 30-day readmission task predicts if a patient will be readmitted in 30 days (00 for no readmission, $1$ for readmission).
Notation Table
Table 1 contains notation symbols and their descriptions used in the paper.
@@ 4.1 Overview
Figure 1: Overall model architecture of our proposed method PRISM. “Rep.” means “Representation”.
@@
  • Feature-Isolated Embedding Module applies GRU and MLP backbone separately to dynamic features and static features. Each dynamic feature learns historical representations over multiple time steps. To align with the original attribute information of each feature, the features are learned in isolation from each other.
@@
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies the graph neural network to learn prototype patients.
@@ 4.2 Feature-Isolated Embedding
In this module, static and dynamic features are learned individually via Multilayer Perceptron (MLP) and Gated Recurrent Unit (GRU) networks, yielding feature representations of unified dimensions $f$.
Static Features Embedding Static features remain constant at each visit. Hence, we opt for a single-layer MLP for simplicity to map each static feature into the feature dimensions $f$:
@@
Dynamic Features Embedding To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling [5]:
@@ 4.3 Feature Confidence Learner
Existing models adopt various ways to enhance patient representations to mitigate the noise introduced by processing sparse EHR data. However, these models often use imputed data and ignore the impact of feature missing status, thus reducing the credibility of the learned hidden representation. We design a measurement called “feature confidence”, which represents the reliability of the input feature values for each patient and each visit. In addition, we have incorporated this measure into the self-attention mechanism as a recalibration module to elevate low-confidence features’ attention.
@@ 4.4 Confidence-Aware Prototype Patient Learner
To comprehensively account for the implications of absent features in the evaluation and identification of analogous patients, we integrate the feature confidence status into our confidence-aware module for discovering cohorts of similar patients. Inspired by GRASP [2]’s graph convolutional network (GCN) framework in EHR modeling, we further compute and integrate the similarity score within the GCN by factoring in the acquired feature confidence pertaining to the individual patient, aiming to facilitate the preferential selection of patients who are more similar to the subject patient.
@@
Prototype Patients Cohort Discovery To compute the enhanced representation information of similar patients, we design the prototype similar patient cohort discovery module by incorporating GCN’s capability to learn relationships between graph nodes. Initially, we utilize the K-Means clustering algorithm to cluster raw patient representation ($F$ recorded or imputed features) into $K$ groups. Subsequently, we identify the center vectors of $K$ clusters as the initial $K$ prototypes $\bm{\mathcal{G}_{k}},k=1,2,\cdots,K$. These prototypes vectors, along with the patient groups, form the nodes of a graph. We then calculate the edge weights for this graph using a predefined similarity measure, resulting in an adjacency matrix $\bm{A}$. As GCN learns across epochs, the graph structure dynamically evolves. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. During the training phase, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adaptively adjusted. Consequently, the most similar and representative prototypes $\bm{\mathcal{G}}$ are identified, and samples within the clusters become more similar to each other. The process is illustrated as: